# Conditional Distribution Compression via the Kernel Conditional Mean Embedding

### v3

@@ B.9 Derivation of ACKH Objective and Gradients
$\nabla_{\bm{x}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})$ $\underset{\text{(a)}}{=}\frac{1}{n}\nabla_{\bm{x}}\text{Tr}\left(\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\nabla_{\bm{x}}\text{Tr}\left(\bar{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\underset{\text{(b)}}{=}\frac{1}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;+\frac{1}{n}\text{Tr}\left(\bar{K}(\bm{x})\nabla_{\bm{x}}\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;+\frac{1}{n}\text{Tr}\left(\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\nabla_{\bm{x}}\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;+\frac{1}{n}\text{Tr}\left(\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\nabla_{\bm{x}}\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\nabla_{\bm{x}}\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\tilde{W}(\bm{x})\nabla_{\bm{x}}\bar{K}(\bm{x})^{T}\right)$
$\underset{\text{(c)}}{=}\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\bar{K}(\bm{x})\nabla_{\bm{x}}\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\nabla_{\bm{x}}\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\tilde{W}(\bm{x})\nabla_{\bm{x}}\bar{K}(\bm{x})^{T}\right),$
@@
$\nabla_{\bm{x}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})$ $=\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{K}(\bm{x})\tilde{W}(\bm{x})\nabla_{\bm{x}}\tilde{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\tilde{W}(\bm{x})\nabla_{\bm{x}}\tilde{K}(\bm{x})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\tilde{W}(\bm{x})\nabla_{\bm{x}}\bar{K}(\bm{x})^{T}\right).$
@@
$A:=\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\in\mathbb{R}^{m\times n},\;\;B:=\bar{L}(\bm{y})\tilde{W}(\bm{x})\in\mathbb{R}^{n\times m},$
$C:=\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{T}\in\mathbb{R}^{m\times n},$
@@
$\nabla_{\bm{x}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})$ $=\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})C\right)-\frac{2}{n}\text{Tr}\left(A^{T}\nabla_{\bm{x}}\tilde{K}(\bm{x})C\right)$
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(B\nabla_{\bm{x}}\tilde{K}(\bm{x})A\right)\ -\frac{2}{n}\text{Tr}\left(B\nabla_{\bm{x}}\bar{K}(\bm{x})^{T}\right)$
$\underset{\text{(a)}}{=}\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})C\right)-\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\tilde{K}(\bm{x})CA^{T}\right)$
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\tilde{K}(\bm{x})AB\right)\ -\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})^{T}B\right)$
where (a) follows from the cyclic property of the trace and the symmetry of the kernels. Now, the cost of computing $A$, $B$ and $C$ is $\mathcal{O}(nm^{2}+m^{3})$, and given these matrices, the cost of computing $D:=CA^{T}\in\mathbb{R}^{m\times m}$ and $E:=AB\in\mathbb{R}^{m\times m}$ is $\mathcal{O}(nm^{2})$. So, we are now in a position where we have to compute
@@
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\tilde{K}(\bm{x})E\right)\ -\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})^{T}B\right).$
@@
where the last equality follows by the symmetry of the kernels. Hence this term, given $D$, can be computed with cost $\mathcal{O}(m)$. Similar derivations hold for $\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})^{T}B\right)$ and $\text{Tr}\left(\nabla_{\bm{x}}\tilde{K}(\bm{x})E\right)$.
@@ B.10 Derivation of ACKIP Objective and Gradients
In ACKIP, we solve the optimisation problem
$\underset{(\tilde{\bm{X}},\tilde{\bm{Y}})\subset\mathcal{X}\times\mathcal{Y}}{\arg\min}\;\;\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\|\tilde{\mu}^{(\tilde{\bm{X}},\tilde{\bm{Y}})}_{Y\mid X=\bm{x}^{\prime}}\right\|^{2}_{\mathcal{H}_{l}}\right]-2\mathbb{E}_{(\bm{x}^{\prime},\bm{y}^{\prime})\sim\mathbb{P}_{X,Y}}\left[\tilde{\mu}^{(\tilde{\bm{X}},\tilde{\bm{Y}})}_{Y\mid X=\bm{x}^{\prime}}(\bm{y}^{\prime})\right].$ (27)
@@
$L_{\tilde{\bm{Y}}\tilde{\bm{Y}}},K_{\tilde{\bm{X}}\tilde{\bm{X}}},W_{\tilde{\bm{X}}\tilde{\bm{X}}}\in\mathbb{R}^{m\times m},\;\;\text{and}\;\;K_{\tilde{\bm{X}}\bm{X}},L_{\bm{Y}\tilde{\bm{Y}}}^{T}\in\mathbb{R}^{m\times n}.$
@@
$\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\tilde{\bm{X}}}\in\mathbb{R}^{m\times m\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\tilde{\bm{X}}}\right]_{ijq}:=\frac{\partial k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})}{\partial\tilde{\bm{x}}_{q}}\in\mathbb{R}^{d},$
$\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\bm{X}}\in\mathbb{R}^{m\times n\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\bm{X}}\right]_{ijq}:=\frac{\partial k(\tilde{\bm{x}}_{i},\bm{x}_{j})}{\partial\tilde{\bm{x}}_{q}}\in\mathbb{R}^{d},$
$\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\in\mathbb{R}^{m\times m\times m\times p},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\right]_{ijq}:=\frac{\partial l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})}{\partial\tilde{\bm{y}}_{q}}\in\mathbb{R}^{p},$
$\nabla_{\tilde{\bm{Y}}}K_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\in\mathbb{R}^{m\times n\times m\times p},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\right]_{ijq}:=\frac{\partial l(\tilde{\bm{y}}_{i},\bm{y}_{j})}{\partial\tilde{\bm{y}}_{q}}\in\mathbb{R}^{p},$
@@
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(B\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}A^{T}\right)-\frac{2}{n}\text{Tr}\left(B^{T}\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\bm{X}}\right)$
@@
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}A^{T}B\right)-\frac{2}{n}\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\bm{X}}B^{T}\right),$
where (a) follows from the cyclic property of the trace and the symmetry of the kernels. Now, the cost of computing $A$, $B$ and $C$ is $\mathcal{O}(nm^{2}+m^{3})$, and given these matrices, the cost of computing $D:=CB\in\mathbb{R}^{m\times m}$ and $E:=A^{T}B\in\mathbb{R}^{m\times m}$ is $\mathcal{O}(nm^{2})$. So, we are now in a position where we have to compute
@@
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}E\right)-\frac{2}{n}\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\bm{X}}B^{T}\right),$
but, following the exact same logic as in section B.4.2, we can compute these traces as sums of products where the majority of the elements of $\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}$ and $\nabla_{\tilde{\bm{X}}}K_{\bm{X}\tilde{\bm{X}}}$ are zeros, and hence much wasted computation and storage can be avoided. Therefore, similar to equations (B.4.2) and (B.4.2), we can compute the quantities $\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\bm{X}\tilde{\bm{X}}}C\right)$ and $\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\bm{X}}B^{T}\right)$ with $\mathcal{O}(mn)$ time and storage cost, and $\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}E\right)$ and $\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}D\right)$ with $\mathcal{O}(m^{2})$ time and storage cost.

### v4

@@ B.9 Derivation of ACKH Objective and Gradients
$\nabla_{\bm{x}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})$ $\underset{\text{(a)}}{=}\frac{1}{n}\nabla_{\bm{x}}\text{Tr}\left(\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\nabla_{\bm{x}}\text{Tr}\left(\bar{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\underset{\text{(b)}}{=}\frac{1}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;+\frac{1}{n}\text{Tr}\left(\bar{K}(\bm{x})\nabla_{\bm{x}}\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;+\frac{1}{n}\text{Tr}\left(\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\nabla_{\bm{x}}\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;+\frac{1}{n}\text{Tr}\left(\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\nabla_{\bm{x}}\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\nabla_{\bm{x}}\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\tilde{W}(\bm{x})\nabla_{\bm{x}}\bar{K}(\bm{x})^{\top}\right)$
$\underset{\text{(c)}}{=}\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\bar{K}(\bm{x})\nabla_{\bm{x}}\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\nabla_{\bm{x}}\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\tilde{W}(\bm{x})\nabla_{\bm{x}}\bar{K}(\bm{x})^{\top}\right),$
@@
$\nabla_{\bm{x}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})$ $=\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{K}(\bm{x})\tilde{W}(\bm{x})\nabla_{\bm{x}}\tilde{K}(\bm{x})\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\tilde{W}(\bm{x})\nabla_{\bm{x}}\tilde{K}(\bm{x})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}(\bm{y})\tilde{W}(\bm{x})\nabla_{\bm{x}}\bar{K}(\bm{x})^{\top}\right).$
@@
$A:=\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\in\mathbb{R}^{m\times n},\;\;B:=\bar{L}(\bm{y})\tilde{W}(\bm{x})\in\mathbb{R}^{n\times m},$
$C:=\tilde{W}(\bm{x})\tilde{L}(\bm{y})\tilde{W}(\bm{x})\bar{K}(\bm{x})^{\top}\in\mathbb{R}^{m\times n},$
@@
$\nabla_{\bm{x}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})$ $=\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})C\right)-\frac{2}{n}\text{Tr}\left(A^{\top}\nabla_{\bm{x}}\tilde{K}(\bm{x})C\right)$
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(B\nabla_{\bm{x}}\tilde{K}(\bm{x})A\right)\ -\frac{2}{n}\text{Tr}\left(B\nabla_{\bm{x}}\bar{K}(\bm{x})^{\top}\right)$
$\underset{\text{(a)}}{=}\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})C\right)-\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\tilde{K}(\bm{x})CA^{\top}\right)$
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\tilde{K}(\bm{x})AB\right)\ -\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})^{\top}B\right)$
where (a) follows from the cyclic property of the trace and the symmetry of the kernels. Now, the cost of computing $A$, $B$ and $C$ is $\mathcal{O}(nm^{2}+m^{3})$, and given these matrices, the cost of computing $D:=CA^{\top}\in\mathbb{R}^{m\times m}$ and $E:=AB\in\mathbb{R}^{m\times m}$ is $\mathcal{O}(nm^{2})$. So, we are now in a position where we have to compute
@@
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\tilde{K}(\bm{x})E\right)\ -\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})^{\top}B\right).$
@@
where the last equality follows by the symmetry of the kernels. Hence this term, given $D$, can be computed with cost $\mathcal{O}(m)$. Similar derivations hold for $\text{Tr}\left(\nabla_{\bm{x}}\bar{K}(\bm{x})^{\top}B\right)$ and $\text{Tr}\left(\nabla_{\bm{x}}\tilde{K}(\bm{x})E\right)$.
@@ B.10 Derivation of ACKIP Objective and Gradients
[…]
$\underset{\tilde{\bm{X}},\tilde{\bm{Y}}}{\arg\min}\;\;\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\|\tilde{\mu}^{(\tilde{\bm{X}},\tilde{\bm{Y}})}_{Y\mid X=\bm{x}^{\prime}}\right\|^{2}_{\mathcal{H}_{l}}\right]-2\mathbb{E}_{(\bm{x}^{\prime},\bm{y}^{\prime})\sim\mathbb{P}_{X,Y}}\left[\tilde{\mu}^{(\tilde{\bm{X}},\tilde{\bm{Y}})}_{Y\mid X=\bm{x}^{\prime}}(\bm{y}^{\prime})\right].$ (27)
@@
$L_{\tilde{\bm{Y}}\tilde{\bm{Y}}},K_{\tilde{\bm{X}}\tilde{\bm{X}}},W_{\tilde{\bm{X}}\tilde{\bm{X}}}\in\mathbb{R}^{m\times m},\;\;\text{and}\;\;K_{\tilde{\bm{X}}\bm{X}},L_{\bm{Y}\tilde{\bm{Y}}}^{\top}\in\mathbb{R}^{m\times n}.$
@@
$\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\tilde{\bm{X}}}\in\mathbb{R}^{m\times m\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\tilde{\bm{X}}}\right]_{ijq}:=\nabla_{\tilde{\bm{x}}_{q}}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})\in\mathbb{R}^{d},$
$\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\bm{X}}\in\mathbb{R}^{m\times n\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}},\bm{X}}\right]_{ijq}:=\nabla_{\tilde{\bm{x}}_{q}}k(\tilde{\bm{x}}_{i},\bm{x}_{j})\in\mathbb{R}^{d},$
$\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\in\mathbb{R}^{m\times m\times m\times p},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\right]_{ijq}:=\nabla_{\tilde{\bm{y}}_{q}}l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})\in\mathbb{R}^{p},$
$\nabla_{\tilde{\bm{Y}}}K_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\in\mathbb{R}^{m\times n\times m\times p},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{Y}}}L_{\tilde{\bm{Y}},\tilde{\bm{Y}}}\right]_{ijq}:=\nabla_{\tilde{\bm{y}}_{q}}l(\tilde{\bm{y}}_{i},\bm{y}_{j})\in\mathbb{R}^{p},$
@@
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(B\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}A^{\top}\right)-\frac{2}{n}\text{Tr}\left(B^{\top}\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\bm{X}}\right)$
@@
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}A^{\top}B\right)-\frac{2}{n}\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\bm{X}}B^{\top}\right),$
where (a) follows from the cyclic property of the trace and the symmetry of the kernels. Now, the cost of computing $A$, $B$ and $C$ is $\mathcal{O}(nm^{2}+m^{3})$, and given these matrices, the cost of computing $D:=CB\in\mathbb{R}^{m\times m}$ and $E:=A^{\top}B\in\mathbb{R}^{m\times m}$ is $\mathcal{O}(nm^{2})$. So, we are now in a position where we have to compute
@@
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}E\right)-\frac{2}{n}\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\bm{X}}B^{\top}\right),$
but, following the exact same logic as in section B.4.2, we can compute these traces as sums of products where the majority of the elements of $\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}$ and $\nabla_{\tilde{\bm{X}}}K_{\bm{X}\tilde{\bm{X}}}$ are zeros, and hence much wasted computation and storage can be avoided. Therefore, similar to equations (B.4.2) and (B.4.2), we can compute the quantities $\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\bm{X}\tilde{\bm{X}}}C\right)$ and $\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\bm{X}}B^{\top}\right)$ with $\mathcal{O}(mn)$ time and storage cost, and $\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}E\right)$ and $\text{Tr}\left(\nabla_{\tilde{\bm{X}}}K_{\tilde{\bm{X}}\tilde{\bm{X}}}D\right)$ with $\mathcal{O}(m^{2})$ time and storage cost.